# references.bib

### v1

@@ 2.2 Structure of spectral extremal graphs
Choose an non-pendant vertex $u$ as the root of $T$ and choose $v\in V(T)$ at maximum distance from $u$. Since $T$ is not a star, $d(v,u)\geq 2$ and also $v$ is a pendant vertex. Let $v\sim_{T}w$. Since $v$ is at maximum distance from $u$ and $w\neq u$, $w$ has a unique neighbor $w^{\prime}$ such that $w^{\prime}$ is not a pendant vertex. Let $T^{\prime}$ be obtained by deleting the edge $wv^{\prime}$ and adding the edge $w^{\prime}v^{\prime}$ for every pendant $v^{\prime}\sim_{T}w^{\prime}$, and note that $T^{\prime}$ has more pendant vertices than $T$. Now $K^{\prime}$ is obtained from $K$ by performing a Kelmans transformation and so $\lambda(K^{\prime})\geq\lambda(K)$ (see e.g. [csikvari2009conjecture]). ∎
@@
Let a longest path in $T$ have endpoints $u$ and $u^{\prime}$ (so that $u,u^{\prime}$ are pendant vertices), and note that if $T$ is not a path then there is a vertex $v$ of degree at least $3$ which is on the path. Let $v\sim w$ with $w\not\in\{u,u^{\prime}\}$. Let $T^{\prime}$ be obtained by deleting $vw$ and adding $uw$ and note that $T^{\prime}$ is a tree of greater diameter than $T$. Let $\mu=\lambda(K)$, with eigenvector $y$, scaled so that the sum of $y_{v}$ over $v$ not in the large part equals $\mu$. For any $x\in V(T)$ one can show $y_{x}=1+\frac{d_{T}(x)}{\mu}+O(n^{-2})$. Therefore
@@
By repeatedly applying 16 to $G_{2}$, it follows that $\lambda(G_{2})\geq\lambda(G_{\mathrm{path}})$. Let $V_{1},\ldots,V_{r}$ be the common partition of $G_{\mathrm{star}}$ and $G_{\mathrm{path}}$, with $v_{i}:=|V_{i}|$. Let $v_{1},\ldots,v_{h}=v_{1}$ and $v_{h+1},\ldots,v_{r}=v_{1}-1$. Let $\mu=\lambda(G_{\mathrm{star}})$, $\rho=\lambda(G_{\mathrm{path}})$, and let $y,z$ be the corresponding eigenvectors, scaled so that $y_{1}:=\frac{1}{\mu}\sum_{v\not\in V_{1}}y_{v}=1$ and $z_{1}:=\frac{1}{\rho}\sum_{v\not\in V_{1}}z_{v}=1$. Similarly to 14, we find $\mu-\rho=O(n^{-1})$. For $v\in V_{1}$, let $d_{s}(v):=d_{G_{\mathrm{star}}[V_{1}]}(v)$ and $d_{p}(v):=d_{G_{\mathrm{path}[V_{1}]}}(v)$, and let $d_{s}^{\prime}(v)$ ($d_{p}^{\prime}(v)$) be the number of walks of length 2 from $v$ in $G_{\mathrm{star}}[V_{1}]$ ($G_{\mathrm{path}}[V_{1}]$). The second-degree eigenvector equations give that for $v\in V_{1}$,
@@
Observe that the number of walks of length two in a path of length $k\geq 4$ is $64k-$ 6. Therefore,
@@
$+(r-h)(v-1)\left(1+\frac{Y_{1}-v_{1}-1}{\mu+v_{1}}+\frac{Y_{1}-v_{1}-1}{(\mu+v_{1})^{2}}\right)+O(n^{-2})$
@@
$+(r-h)(v_{1}-1)\left(1+\frac{Z_{1}-v_{1}-1}{\mu+v_{1}}+\frac{Z_{1}-v_{1}-1}{(\mu+v_{1})^{2}}\right)+O(n^{-2}).$
@@
Suppose not, and assume without loss of generality that $w_{1}=w_{2}+2$. Let $w_{2},\ldots,w_{h}=w_{1}-2$ and $w_{h+1},\ldots,w_{r}=w_{1}-1$. Let $G^{\prime}$ be obtained from $G$ by converting every vertex in $W^{m}$ to an ordinary vertex, and let $a=e(G^{\prime})-e(G)=O(1)$. Let $G^{\prime\prime}$ be obtained from $G^{\prime}$ by deleting a pendant edge in $G^{\prime}[W_{1}]$. Then $e(G^{\prime\prime})=e(G)-a-1$, and $G^{\prime\prime}$ has an isolated vertex $u^{\prime}$ in $G^{\prime\prime}[W_{1}]$. Let $\mu=\lambda(G^{\prime\prime})$, $\rho=\lambda(G_{12}^{\prime\prime})$, where $G_{12}^{\prime\prime}$ is obtained by transferring the isolated vertex $u^{\prime}$ to $W_{2}$; then $G_{12}^{\prime\prime}$ is $\mathcal{F}$-free and $e(G_{12}^{\prime\prime})=e(G^{\prime\prime})+1=e(G)-a$. Let $y$ be the principal eigenvector of $G^{\prime\prime}$ and $z$ be the principal eigenvector of $G_{12}^{\prime\prime}$, scaled so that $y_{1}:=\frac{1}{\mu}\sum_{v\not\in W_{1}}y_{v}=1$ and $z_{1}:=\frac{1}{\rho}\sum_{v\not\in W_{1}-\{u^{\prime}\}}z_{v}=1$. Similarly to 14, we find $\rho-\mu=O(n^{-1})$.
@@
$Y_{1}$ $=w_{1}+\frac{2e(W_{1})}{\mu}+\frac{\sum_{v\in W_{1}-\{u^{\prime}\}}d_{W_{1}}^{\prime}(v)}{\mu^{2}}+O(n^{-2})$
@@
We now obtain an initial estimate on $\mu$. We get
@@
$=$  $\,1-\frac{rw_{2}}{\mu+w_{2}}+\frac{Y_{1}-w_{2}-r+h-2}{\mu+w_{2}}-\frac{w_{2}(Y_{1}-w_{2}-r+h-2)}{(\mu+w_{2})^{2}}+O(n^{-1}).$
@@
Note that $G_{12}^{\prime\prime}$ is obtained from $T_{n,r}$ by embedding trees of order $k$ in a large part. By 1 and applying a vertex transfer if necessary, some graph $H\in\mathrm{EX}(n,\mathcal{F})$ is obtained from $T_{n,r}$ by embedding trees of order $k$ in a large part, then making a bounded number of edge modifications. There is a graph $H^{\prime}$ obtained from $H$ by a bounded number of edge modifications and a graph $G_{12}^{\prime\prime\prime}$ obtained from $G_{12}^{\prime\prime}$ by a bounded number of edge modifications, so that each graph is obtained from $T_{n,r}$ by embedding the same number of trees of order $k$ in the large part; let $b=e(H)-e(H^{\prime})=O(1)$, $a^{\prime}=e(G_{12}^{\prime\prime\prime})-e(G_{12}^{\prime\prime})=O(1).$ Now if $G\in\mathrm{EX}(n,\mathcal{F})$ we are done, so we may assume $b-a^{\prime}-a=e(H)-e(G)\geq 1$.
@@
$\lambda(H)-\lambda(G)=$ $\,\lambda(H)-\lambda(H^{\prime})+\lambda(H^{\prime})-\lambda(G_{21}^{\prime\prime\prime})+\lambda(G_{21}^{\prime\prime\prime})-\lambda(G_{21}^{\prime\prime})$
$+\lambda(G_{21}^{\prime\prime})-\lambda(G^{\prime\prime})+\lambda(G^{\prime\prime})-\lambda(G)$
$=$ $\,\left[2b+0-2a^{\prime}+2-\frac{2}{r}+\frac{4(k-1)}{kr}-2(a+1)\right]\frac{1}{n}+O(n^{-2})>0$
@@
$\lambda(H)-\lambda(G)=$ $\,\lambda(H)-\lambda(H^{\prime})+\lambda(H^{\prime})-\lambda(G_{21}^{\prime\prime\prime})+\lambda(G_{21}^{\prime\prime\prime})-\lambda(G_{21}^{\prime\prime})$
$+\lambda(G_{21}^{\prime\prime})-\lambda(G^{\prime\prime})+\lambda(G^{\prime\prime})-\lambda(G)$
$\geq$ $\,\left[2b-\frac{k-5+6/k}{r-1}-2a^{\prime}+2-\frac{2}{r}+\frac{4(k-1)}{kr}-2(a+1)\right]\frac{1}{n}+O(n^{-2})$
@@
Case 2: $w_{1}=v_{1}-1.$ Then without loss of generality, $w_{2}=w_{1}+1=v_{1}$. Let $G^{\prime}$ be obtained from $G$ by replacing every vertex in $W^{m}$ with an ordinary vertex. Let $a=e(G)-e(G^{\prime})=O(1)$.
@@
Also $z_{u^{\prime}}=1$. Then
@@
$z_{v}=\begin{cases}1+\frac{Y_{1}-w_{1}}{\mu+w_{1}}-\frac{Y_{1}-w_{1}}{(\mu+w_{1})^{2}}+O(n^{-3}),&3\leq i\leq h\\ 1+\frac{Y_{1}+w_{1}-1}{\mu+w_{1}}+O(n^{-3}),&i=2\text{ or }h+1\leq i\leq r.\end{cases}$
@@
Subcase 2.1: $k\in\{1,2,3\}$. Then $H$ can be obtained from $G_{21}^{\prime}$ by a bounded number of edge modifications. Let $b=e(H)-e(G_{21}^{\prime})=e(H)-e(G^{\prime})>e(G)=a$. Then by 41 and 18, we have
@@
$\geq\left[2b-\frac{4(k-1)}{kr}-2a\right]\frac{1}{n}+O(n^{-2}).$
The bracketed expression is positive, so this is a contradiction.
Subcase 2.2: $k\geq 4$. Let $H^{\prime}$ be obtained from $H$ by deleting some $b=O(1)$ edges, and $G_{21}^{\prime\prime}$ be obtained from $G_{21}^{\prime}$ by deleting some $a^{\prime}$ edges, so that $H^{\prime}$ and $G_{21}^{\prime\prime}$ are each obtained by embedding the same number of trees of order $k$ in a large part of $T_{n,r}$. Note that $e(H)>e(G)$ gives $b-a-a^{\prime}>0$. Then by 41, 43, and 18, we have
@@
$\geq\left[2b-\frac{k-5+6/k}{r-1}-2a^{\prime}-\frac{4(k-1)}{kr}-2a\right]\frac{1}{n}+O(n^{-2})$
@@
Observe that in Subcase 2.2 of the proof of Theorem 2, if $e(H)=e(G)+1$ and all the trees embedded in $H$ are paths while all the trees embedded in $G$ are stars, then according to 43, the equality
@@
holds. In the next section we will show that there is a forbidden family $\mathcal{F}$ such that some $H\in\mathrm{EX}(n,\mathcal{F})$ and some $\mathcal{F}$-free graph $G$ in fact have such a form, and use this to prove 1.
@@ 3.1 Large extremal number
Suppose that $r\geq 3$, $n\not\equiv 0\pmod{r}$, and $\lfloor n/r\rfloor\equiv 0\pmod{k}$ (note there are infinitely many such $n$). Let $G$ be obtained by embedding $U(K_{1,k-1},\lfloor n/r\rfloor)$ in a small part of $T_{n,r}$ and let $H$ be obtained by embedding $U(P_{k},\lceil n/r\rceil)$ in a large part of $T_{n,r}$ and then extending one of the paths by a single edge, so that $e(H)=e(G)+1$. Let $V_{1}\sqcup\cdots\sqcup V_{r}$ and $U_{1}\sqcup\cdots\sqcup U_{r}$ be the corresponding partitions of $V(G)$ and $V(H)$, respectively, so that the trees are embedded in $V1$ and $U_{1}$. Since
@@
By the observation below the proof of Theorem 2, this implies that $\lambda(H)<\lambda(G)$. Thus, to prove 1 is suffices to construct a family $\mathcal{F}$ such that $\mathrm{EX}(n,\mathcal{F})=\{H\}$ and $G$ is $\mathcal{F}$-free. Let $m\geq\max\{2,k-1\}$ and let $\mathcal{F}$ consist of the following graphs.
@@
    $T_{k}+T_{m(r-1),r-1}$ for any tree $T_{k}$ on $k$ vertices for which $T_{k}\not\simeq K_{1,k-1},P_{k}$
@@
    $T_{k+1}+T_{m(r-1),r-1}$ for any tree $T_{k+1}$ on $k+1$ vertices for which $T_{k+1}\not\simeq P_{k+1}$
@@
Consider any copy $K$ of $T_{m(r-1),r-1}$ which appears in $H$ and assume for a contradiction that a forbidden subgraph of the form $F+K$ appears. If $V(K)\subseteq V-U_{1}$, then $N(V(K))=U_{1}$, so no forbidden graph of the form $F+K$ appears, by the definition of of $H[U_{1}]$. Otherwise, $V(K)\ni v\in U_{1}$. If $|V(K)\cap U_{1}|\geq 3$, then $N(V(K)\cap U_{1})\cap U_{1}=\emptyset$, hence $V(F)\cap U_{1}=\emptyset$. Then $V(F)$ intersects two different parts $U_{2},U_{3}$, implying that $K$ is contained in the $r-2$ parts $U_{1},U_{4},\ldots,U_{r}$. But one can check $|V(K)\cap U_{i}|\leq m$ since $\Delta(H[U_{i}])\leq 2\leq m$ so $|V(K)|\leq(r-2)m$, a contradiction. Thus, $|V(K)\cap U_{1}|\leq 2$, so $|V(K)-U_{1}|\geq m(r-1)-2$, and thus $V(K)\cap U_{i}\neq\emptyset$ for each $i$. Hence, $N(V(K))\subseteq U_{1}$, and the previous argument applies. This covers forbidden subgraphs of forms (1) through (5). Finally, if $H$ contains a copy $K$ of $K_{r+2}$, then $|V(K)\cap U_{i}|\leq 1$ for each $i>1$ and $|V(K)\cap U_{1}|\leq 2$. Hence $|V(K)|\leq r+1$, a contradiction. ∎
@@
Consider any copy $K$ of $T_{m(r-1),r-1}$ which appears in $G$. If $V(K)\subseteq V-V_{1}$, then $N(V(K))=V_{1}$, so no forbidden graph of the form $F+K$ appears, by the definition of $G[V_{1}]$. Otherwise, $V(K)\ni v\in V_{1}$. If $|V(K)\cap V_{1}|\geq k$, then $N(V(K)\cap V_{1})\cap V_{1}=\emptyset$, hence $V(F)\cap V_{1}=\emptyset$. Then $V(F)$ intersects two different parts $V_{2},V_{3}$, implying that $K$ is contained in the $r-2$ parts $V_{1},V_{4},\ldots,V_{r}$. But one can check $|V(K)\cap V_{i}|\leq m$ since $\Delta(G[V_{i}])\leq k-1\leq m$ so $|V(K)|\leq(r-2)m$, a contradiction. Thus, $|V(K)\cap V_{1}|\leq k-1$, so $|V(K)-V_{1}|\geq m(r-1)-k+1$, and thus $V(K)\cap V_{i}\neq\emptyset$ for each $i$ since $m\geq k-1$. Hence $N(V(K))\subseteq V_{1}$, and the previous argument applies. This covers forbidden subgraphs of forms (1) through (5). Finally, if $G$ contains a copy $K$ of $K_{r+2}$, then $|V(K)\cap U_{i}|\leq 1$ for each $i>1$ and $|V(K)\cap V_{1}|\leq 2$. Hence $|V(K)|\leq r+1$, a contradiction. ∎

### v2

@@ 2.2 Structure of spectral extremal graphs
Choose a non-pendant vertex $u$ as the root of $T$ and choose $v\in V(T)$ at maximum distance from $u$. Since $T$ is not a star, $d(v,u)\geq 2$ and also $v$ is a pendant vertex. Let $v\sim_{T}w$. Since $v$ is at maximum distance from $u$ and $w\neq u$, $w$ has a unique neighbor $w^{\prime}$ such that $w^{\prime}$ is not a pendant vertex. Let $T^{\prime}$ be obtained by deleting the edge $wv^{\prime}$ and adding the edge $w^{\prime}v^{\prime}$ for every pendant $v^{\prime}\sim_{T}w^{\prime}$, and note that $T^{\prime}$ has more pendant vertices than $T$. Now $K^{\prime}$ is obtained from $K$ by performing a Kelmans transformation and so $\lambda(K^{\prime})\geq\lambda(K)$ (see e.g. [csikvari2009conjecture]). ∎
@@
Let a longest path $P$ in $T$ have endpoints $u$ and $u^{\prime}$ (so that $u,u^{\prime}$ are pendant in $T$), and note that if $T$ is not a path then there is a vertex $v$ of degree at least $3$ which is on $P$. Let $v\sim w$ with $w\not\in\{u,u^{\prime}\}$. Let $T^{\prime}$ be obtained from $T$ by deleting $vw$ and adding $uw$ and note that $T^{\prime}$ is a tree of greater diameter than $T$. Let $\mu=\lambda(K)$, with eigenvector $y$, scaled so that the sum of $y_{v}$ over $v$ not in the large part equals $\mu$. For any $x\in V(T)$ one can show $y_{x}=1+\frac{d_{T}(x)}{\mu}+O(n^{-2})$. Therefore
@@
By repeatedly applying 16 to $G_{2}$, it follows that $\lambda(G_{2})\geq\lambda(G_{\mathrm{path}})$. Let $V_{1},\ldots,V_{r}$ be the common partition of $G_{\mathrm{star}}$ and $G_{\mathrm{path}}$, with $v_{i}:=|V_{i}|$. Let $v_{1},\ldots,v_{h}=v_{1}$ and $v_{h+1},\ldots,v_{r}=v_{1}-1$. Let $\mu=\lambda(G_{\mathrm{star}})$, $\rho=\lambda(G_{\mathrm{path}})$, and let $y,z$ be the corresponding eigenvectors, scaled so that $y_{1}:=\frac{1}{\mu}\sum_{v\not\in V_{1}}y_{v}=1$ and $z_{1}:=\frac{1}{\rho}\sum_{v\not\in V_{1}}z_{v}=1$. Similarly to 14, we find $\mu-\rho=O(n^{-1})$. For $v\in V_{1}$, let $d_{s}(v):=d_{G_{\mathrm{star}}[V_{1}]}(v)$ and $d_{p}(v):=d_{G_{\mathrm{path}[V_{1}]}}(v)$, and let $d_{s}^{\prime}(v)$ ($d_{p}^{\prime}(v)$) be the number of walks of length 2 starting from $v$ in $G_{\mathrm{star}}[V_{1}]$ ($G_{\mathrm{path}}[V_{1}]$). The second-degree eigenvector equations give that for $v\in V_{1}$,
@@
Observe that the number of walks of length two in a path of order $k\geq 4$ is $64k-$ 6 and the number of walks of length two in $K_{1,k-1}$ is $k(k-1)$. Therefore,
@@
$+(r-h)(v_{1}-1)\left(1+\frac{Y_{1}-v_{1}+1}{\mu+v_{1}}+\frac{Y_{1}-v_{1}+1}{(\mu+v_{1})^{2}}\right)+O(n^{-2})$
@@
$+(r-h)(v_{1}-1)\left(1+\frac{Z_{1}-v_{1}+1}{\mu+v_{1}}+\frac{Z_{1}-v_{1}+1}{(\mu+v_{1})^{2}}\right)+O(n^{-2}).$
@@
Suppose not, and assume without loss of generality that $w_{1}=w_{2}+2$. Let $w_{2},\ldots,w_{h}=w_{1}-2$ and $w_{h+1},\ldots,w_{r}=w_{1}-1$. Let $G^{\prime}$ be obtained from $G$ by converting every vertex in $W^{m}$ to a $W$-ordinary vertex, and let $a=e(G^{\prime})-e(G)=O(1)$. Let $G^{\prime\prime}$ be obtained from $G^{\prime}$ by deleting a pendant edge in $G^{\prime}[W_{1}]$. Then $e(G^{\prime\prime})=e(G)+a-1$, and $G^{\prime\prime}$ has an isolated vertex $u^{\prime}$ in $G^{\prime\prime}[W_{1}]$. Let $\mu=\lambda(G^{\prime\prime})$, $\rho=\lambda(G_{12}^{\prime\prime})$, where $G_{12}^{\prime\prime}$ is obtained by transferring $u^{\prime}$ to $W_{2}$; then $G_{12}^{\prime\prime}$ is $\mathcal{F}$-free and $e(G_{12}^{\prime\prime})=e(G^{\prime\prime})+1=e(G)+a$. Let $y$ be the principal eigenvector of $G^{\prime\prime}$ and $z$ be the principal eigenvector of $G_{12}^{\prime\prime}$, scaled so that $y_{1}:=\frac{1}{\mu}\sum_{v\not\in W_{1}}y_{v}=1$ and $z_{1}:=\frac{1}{\rho}\sum_{v\not\in W_{1}-\{u^{\prime}\}}z_{v}=1$. Similarly to 14, we find $\rho-\mu=O(n^{-1})$.
@@
$Y_{1}$ $=w_{1}+\frac{2(e_{G}(W_{1})-1)}{\mu}+\frac{\sum_{v\in W_{1}-\{u^{\prime}\}}d_{W_{1}}^{\prime}(v)}{\mu^{2}}+O(n^{-2})$
@@
We now obtain an initial estimate on $\mu$. We have
@@
$=$  $\,1-\frac{rw_{2}}{\mu+w_{2}}+\frac{Y_{1}-w_{2}-r+h-2}{\mu+w_{2}}-\frac{w_{2}(Y_{1}-w_{2}-r+h-2)}{(\mu+w_{2})^{2}}+O(n^{-2}).$
@@
Note that $G_{12}^{\prime\prime}$ is obtained from $T_{n,r}$ by embedding trees of order $k$ in a large part, with possible 1 smaller tree. By 1 and applying a vertex transfer if necessary, some graph $H\in\mathrm{EX}(n,\mathcal{F})$ is obtained from $T_{n,r}$ by embedding trees of order $k$ in a large part, then making a bounded number of edge modifications. There is a graph $H^{\prime}$ obtained from $H$ by a bounded number of edge modifications and a graph $G_{12}^{\prime\prime\prime}$ obtained from $G_{12}^{\prime\prime}$ by a bounded number of edge modifications, so that each graph is obtained from $T_{n,r}$ by embedding the same number of trees of order $k$ in the large part; let $b=e(H)-e(H^{\prime})=O(1)$, $a^{\prime}=e(G_{12}^{\prime\prime\prime})-e(G_{12}^{\prime\prime})=O(1).$ Now if $G\in\mathrm{EX}(n,\mathcal{F})$ we are done, so we may assume $b+a^{\prime}+a=e(H)-e(G)\geq 1$.
@@
$\lambda(H)-\lambda(G)=$ $\,\lambda(H)-\lambda(H^{\prime})+\lambda(H^{\prime})-\lambda(G_{12}^{\prime\prime\prime})+\lambda(G_{12}^{\prime\prime\prime})-\lambda(G_{12}^{\prime\prime})$
$+\lambda(G_{12}^{\prime\prime})-\lambda(G^{\prime\prime})+\lambda(G^{\prime\prime})-\lambda(G)$
$=$ $\,\left[2b+0+2a^{\prime}+2-\frac{2}{r}+\frac{4(k-1)}{kr}+2(a-1)\right]\frac{1}{n}+O(n^{-2})>0$
@@
$\lambda(H)-\lambda(G)=$ $\,\lambda(H)-\lambda(H^{\prime})+\lambda(H^{\prime})-\lambda(G_{12}^{\prime\prime\prime})+\lambda(G_{12}^{\prime\prime\prime})-\lambda(G_{12}^{\prime\prime})$
$+\lambda(G_{12}^{\prime\prime})-\lambda(G^{\prime\prime})+\lambda(G^{\prime\prime})-\lambda(G)$
$\geq$ $\,\left[2b-\frac{k-5+6/k}{r-1}+2a^{\prime}+2-\frac{2}{r}+\frac{4(k-1)}{kr}+2(a-1)\right]\frac{1}{n}+O(n^{-2})$
@@
Case 2: $w_{1}=v_{1}-1.$ Then without loss of generality, $w_{2}=w_{1}+1=v_{1}$. Let $G^{\prime}$ be obtained from $G$ by replacing every vertex in $W^{m}$ with an ordinary vertex. Let $a=e(G^{\prime})-e(G)=O(1)$.
@@
Also $z_{u^{\prime}}=1$. Let $Y_{1}=\sum_{v\in W_{1}}y_{v},$ $Z_{1}=\sum_{W_{1}\cup\{u^{\prime}\}}z_{v}$. Then
@@
$z_{v}=\begin{cases}1+\frac{Y_{1}-w_{1}}{\mu+w_{1}}-\frac{Y_{1}-w_{1}}{(\mu+w_{1})^{2}}+O(n^{-3}),&3\leq i\leq h\\ 1+\frac{Y_{1}-w_{1}+1}{\mu+w_{1}}+O(n^{-3}),&i=2\text{ or }h+1\leq i\leq r.\end{cases}$
@@
Subcase 2.1: $k\in\{1,2,3\}$. Then $H$ can be obtained from $G_{21}^{\prime}$ by a bounded number of edge modifications. Let $b=e(H)-e(G_{21}^{\prime})=e(H)-e(G^{\prime})=e(H)-e(G)-a$, so $b+a\geq 1$. Then by 41 and 18, we have
@@
$\geq\left[2b-\frac{4(k-1)}{kr}+2a\right]\frac{1}{n}+O(n^{-2})>0,$
a contradiction.
Subcase 2.2: $k\geq 4$. Let $H^{\prime}$ be obtained from $H$ by deleting some $b=O(1)$ edges, and $G_{21}^{\prime\prime}$ be obtained from $G_{21}^{\prime}$ by deleting some $a^{\prime}=O(1)$ edges, so that $H^{\prime}$ and $G_{21}^{\prime\prime}$ are each obtained by embedding the same number of trees of order $k$ in a large part of $T_{n,r}$. Note that $e(H)>e(G)$ gives $b+a-a^{\prime}\geq 1$. Then by 41, 43, and 18, we have
@@
$\geq\left[2b-\frac{k-5+6/k}{r-1}-2a^{\prime}-\frac{4(k-1)}{kr}+2a\right]\frac{1}{n}+O(n^{-2})$
@@
Observe that in Subcase 2.2 of the proof of Theorem 2, if $e(H)=e(G)+1$, $k\geq 4$, and all the trees embedded in $H$ are paths while all the trees embedded in $G$ are stars, then according to 43, the equality
@@
holds. In the next section we will show that there is a forbidden family $\mathcal{F}$ such that some $H\in\mathrm{EX}(n,\mathcal{F})$ and some $\mathcal{F}$-free graph $G$ in fact satisfy these conditions, and then use this to prove 1.
@@ 3.1 Large extremal number
Suppose that $r\geq 3$, $n\equiv 1\pmod{r}$ is large enough, and $\lfloor n/r\rfloor=(n-1)/r\equiv 0\pmod{k}$ (note there are infinitely many such $n$). Let $G$ be obtained by embedding $U(K_{1,k-1},\lfloor n/r\rfloor)$ in a small part of $T_{n,r}$ and let $H$ be obtained by embedding $U(P_{k},\lceil n/r\rceil)$ in a large part of $T_{n,r}$ and then extending one of the paths by a single edge within the part, so that $e(H)=e(G)+1$. Let $V_{1}\sqcup\cdots\sqcup V_{r}$ and $U_{1}\sqcup\cdots\sqcup U_{r}$ be the corresponding partitions of $V(G)$ and $V(H)$, respectively, so that the trees are embedded in $V_{1}$ and $U_{1}$. Since
@@
and in particular $k\geq 4$. Thus, by the observation below the proof of Theorem 2, this implies that $\lambda(H)<\lambda(G)$. So, to prove 1 it suffices to construct a family $\mathcal{F}$ such that $\mathrm{EX}(n,\mathcal{F})=\{H\}$ and $G$ is $\mathcal{F}$-free. Let $m>\max\{4,k-1\}$ and let $\mathcal{F}$ be the union of the following graph families.
@@
    $T_{k}+T_{m(r-1),r-1}$ for every tree $T_{k}$ on $k$ vertices for which $T_{k}\not\simeq K_{1,k-1},P_{k}$
@@
    $T_{k+1}+T_{m(r-1),r-1}$ for every tree $T_{k+1}$ on $k+1$ vertices for which $T_{k+1}\not\simeq P_{k+1}$
@@
Consider any copy $K$ of $T_{m(r-1),r-1}$ which appears in $H$ and assume for a contradiction that a forbidden subgraph of the form $F+K$ appears. If $V(K)\subseteq V-U_{1}$, then $N(V(K))=U_{1}$, so no forbidden graph of the form $F+K$ appears, by the definition of $H[U_{1}]$. Otherwise, $V(K)\ni v\in U_{1}$. If $|V(K)\cap U_{1}|\geq 3$, then $N(V(K)\cap U_{1})\cap U_{1}=\emptyset$, hence $V(F)\cap U_{1}=\emptyset$. Then $V(F)$ intersects two different parts $U_{2},U_{3}$, implying that $K$ is contained in the $r-2$ parts $U_{1},U_{4},\ldots,U_{r}$. But one can check $|V(K)\cap U_{i}|\leq m$ since $\Delta(H[U_{i}])\leq 2<m$ so $|V(K)|\leq(r-2)m$, a contradiction. Thus, $|V(K)\cap U_{1}|\leq 2$, so $|V(K)-U_{1}|\geq m(r-1)-2$, and thus $V(K)\cap U_{i}\neq\emptyset$ for each $i$ as $m(r-2)+2<m(r-1)-2$. Hence, $N(V(K))\subseteq U_{1}$, and the previous argument applies. This covers forbidden subgraphs of forms (1) through (5). Finally, if $H$ contains a copy $K$ of $K_{r+2}$, then $|V(K)\cap U_{i}|\leq 1$ for each $i>1$ and $|V(K)\cap U_{1}|\leq 2$. Hence $|V(K)|\leq r+1$, a contradiction. ∎
@@
Consider any copy $K$ of $T_{m(r-1),r-1}$ which appears in $G$. If $V(K)\subseteq V-V_{1}$, then $N(V(K))=V_{1}$, so no forbidden graph of the form $F+K$ appears, by the definition of $G[V_{1}]$. Otherwise, $V(K)\ni v\in V_{1}$. If $|V(K)\cap V_{1}|\geq k$, then $N(V(K)\cap V_{1})\cap V_{1}=\emptyset$, hence $V(F)\cap V_{1}=\emptyset$. Then $V(F)$ intersects two different parts $V_{2},V_{3}$, implying that $K$ is contained in the $r-2$ parts $V_{1},V_{4},\ldots,V_{r}$. But one can check $|V(K)\cap V_{i}|\leq m$ since $\Delta(G[V_{i}])\leq k-1<m$ so $|V(K)|\leq(r-2)m$, a contradiction. Thus, $|V(K)\cap V_{1}|\leq k-1$, so $|V(K)-V_{1}|\geq m(r-1)-k+1$, and thus $V(K)\cap V_{i}\neq\emptyset$ for each $i$ since $m\geq k-1$. Hence $N(V(K))\subseteq V_{1}$, and the previous argument applies. This covers forbidden subgraphs of forms (1) through (5). Finally, if $G$ contains a copy $K$ of $K_{r+2}$, then $|V(K)\cap U_{i}|\leq 1$ for each $i>1$ and $|V(K)\cap V_{1}|\leq 2$. Hence $|V(K)|\leq r+1$, a contradiction. ∎